# Double Yangians and quantum vertex algebras, I

### v1

@@ 5.2. Universal vacuum $\widehat{\mathcal{DY}}(A)$-module ${\mathcal{V}}_{A}(\ell)$
(5.43) $\tilde{h}_{i}(z)\tilde{h}_{j}(w)-\tilde{h}_{j}(w)\tilde{h}_{i}(z)=[a_{i,j}]_{q^{\partial_{w}}}[\ell]_{q^{\partial_{w}}}((z-w+\ell\hbar)^{-2}-(w-z+\ell\hbar)^{-2}),$
(5.44) $\tilde{h}_{i}(z)\tilde{e}_{j}(w)-\tilde{e}_{j}(w)\tilde{h}_{i}(z)=\tilde{e}_{j}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w+\ell\hbar)^{-1}-(w-z+\ell\hbar)^{-1}),$
(5.45) $\tilde{h}_{i}(z)\tilde{f}_{j}(w)-\tilde{f}_{j}(w)\tilde{h}_{i}(z)=-\tilde{f}_{j}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w+\ell\hbar)^{-1}-(w-z+\ell\hbar)^{-1}),$
(5.46) $(z-w-a_{i,j}\hbar)\tilde{e}_{i}(z)\tilde{e}_{j}(w)=(z-w+a_{i,j}\hbar)\tilde{e}_{j}(w)\tilde{e}_{i}(z),$
(5.47) $(z-w-a_{i,j}\hbar)\tilde{f}_{i}(z)\tilde{f}_{j}(w)=(z-w+a_{i,j}\hbar)\tilde{f}_{j}(w)\tilde{f}_{i}(z),$
(5.48) $(z-w+a_{i,j}\hbar)\tilde{e}_{i}(z)\tilde{f}_{j}(w)=(z-w-a_{i,j}\hbar)\tilde{f}_{j}(w)\tilde{e}_{i}(z)\quad\text{ if }i\neq j,$
(5.49) $(z-w)(z-w+2\ell\hbar)\!\left((z-w+2\hbar)\tilde{e}_{i}(z)\tilde{f}_{i}(w)-(z-w-2\hbar)\tilde{f}_{i}(w)\tilde{e}_{i}(z)\right)\!=0,$
and the order-$2$ Serre relations for $(\tilde{e}_{i}(z_{1}),\tilde{e}_{i}(z_{1}),\tilde{e}_{j}(w))$ and $(\tilde{f}_{i}(z_{1}),\tilde{f}_{i}(z_{1}),\tilde{f}_{j}(w))$ for $i,j\in I$ with $a_{i,j}=-1$.
Note that for any $r,s\in{\mathbb{Z}}$, the coefficients of monomial $z^{r}w^{s}$ in all the relations above indeed lie in $\widetilde{\mathcal{U}}$$\hbar$]].
@@
(5.59) $(\hat{e}_{i})_{0}(\hat{f}_{j})-\frac{1}{2\hbar}\delta_{i,j}({\bf 1}-C_{i}),\ \ (\hat{e}_{i})_{n}(\hat{f}_{j})-\delta_{i,j}\ell(-2\ell\hbar)^{n-1}C_{i}\ \ \text{ for }i,j\in I,\ n\geq 1$
and vectors
(5.60) $(\hat{e}_{i})_{n}(\hat{e}_{j}),\ (\hat{f}_{i})_{n}(\hat{f}_{j})\quad\text{ for }i,j\in I\ \text{with }a_{i,j}=0\ \text{and for }n\geq 0.$
@@
$Y(\check{e}_{i},z)Y(\check{f}_{j},w)\sim\left(\frac{w-z+a_{i,j}\hbar}{w-z-a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z).$
@@
(5.65) $Y(\check{e}_{i},z)Y(\check{f}_{j},w)-\!\left(\frac{w-z+a_{i,j}\hbar}{w-z-a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z)$
@@
(5.66) $(\check{e}_{i})_{0}(\check{f}_{j})=\frac{1}{2\hbar}\delta_{i,j}({\bf 1}-\mathfrak{c}_{i}),\ \ (\check{e}_{i})_{n}(\check{f}_{j})=\delta_{i,j}\ell(-2\ell\hbar)^{n-1}\mathfrak{c}_{i}$
@@
(5.67) $Y(\check{e}_{i},z)Y(\check{f}_{j},w)-\!\left(\frac{w-z+a_{i,j}\hbar}{w-z-a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z)$
@@
Let $i,j\in I$ with $a_{i,j}=0$. We have $(\check{e}_{i})_{n}(\check{e}_{j})=0=(\check{f}_{i})_{n}(\check{f}_{j})$ for $n\geq 0$. With relations (5.46) and (5.47), using the arguments above we get
(5.69) $Y(\check{e}_{i},z)Y(\check{e}_{j},w)=Y(\check{e}_{j},w)Y(\check{e}_{i},z),$
(5.70) $Y(\check{f}_{i},z)Y(\check{f}_{j},w)=Y(\check{f}_{j},w)Y(\check{f}_{i},z).$
@@
$[Y(\check{h}_{i},z),Y(\check{h}_{j},w)]=[a_{i,j}]_{q^{\partial_{w}}}[\ell]_{q^{\partial_{w}}}\left((z-w+\ell\hbar)^{-2}-(w-z+\ell\hbar)^{-2}\right)\!,$
$[Y(\check{h}_{i},z),Y(\check{e}_{j}^{\pm},w)]=\pm Y(\check{e}_{j}^{\pm},w)[a_{i,j}]_{q^{\partial_{w}}}\left((z-w+\ell\hbar)^{-1}+(w-z+\ell\hbar)^{-1}\right)\!,$
$(z-w-a_{i,j}\hbar)Y(\check{e}_{i}^{\pm},z)Y(\check{e}_{j}^{\pm},w)=(z-w+a_{i,j}\hbar)Y(\check{e}_{j}^{\pm},w)Y(\check{e}_{i}^{\pm},z),$
$Y(\check{e}_{i}^{\pm},z)Y(\check{e}_{j}^{\pm},w)=Y(\check{e}_{j}^{\pm},w)Y(\check{e}_{i}^{\pm},z)\quad\text{ if }a_{i,j}=0,$
$Y(\check{e}_{i},z)Y(\check{f}_{j},w)-\left(\frac{w-z+a_{i,j}\hbar}{w-z-a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z)$
$\quad=\delta_{i,j}\frac{1}{2\hbar}\!\left(\!z^{-1}\delta\!\left(\frac{w}{z}\right)-Y\!\left(q^{(1-\ell)\mathcal{D}}E^{-}(-\check{h}_{i},-2\hbar){\bf 1},w\right)z^{-1}\delta\!\left(\frac{w-2\ell\hbar}{z}\right)\!\right)\!,$
where $\mathcal{D}$ denotes the $D$-operator of ${\mathcal{V}}_{A}(\ell)$. In addition, for $i,j\in I$ with $a_{i,j}=-1$, the Serre relations for $(Y(\check{e}_{i}^{\pm},z_{1}),Y(\check{e}_{i}^{\pm},z_{2}),Y(\check{e}_{j}^{\pm},w))$ hold. On the other hand, for any ${\mathcal{V}}_{A}(\ell)$-module $(W,Y_{W})$, all of these relations hold on $W$ by Propositions 3.26 and 3.22 and by Lemma 3.25. Furthermore, we have
@@
(L1) $[h_{i}(z),h_{j}(w)]=a_{i,j}\frac{\partial}{\partial{w}}z^{-1}\delta\!\left(\frac{w}{z}\right){\bf k},$
(L2) $[h_{i}(z),e_{j}^{\pm}(w)]=\pm a_{i,j}e_{j}^{\pm}(w)z^{-1}\delta\!\left(\frac{w}{z}\right)\!,$
(L3) $[e_{i}^{+}(z),e_{j}^{-}(w)]=\delta_{i,j}\left(\!h_{i}(w)z^{-1}\delta\!\left(\frac{w}{z}\right)+{\bf k}\frac{\partial}{\partial{w}}z^{-1}\delta\!\left(\frac{w}{z}\right)\right)\!,$
@@
(L5) $(z-w)[e_{i}^{\pm}(z),e_{j}^{\pm}(w)]=0\quad\textnormal{{if }}a_{i,j}=-1,$
(L6) $[e_{i}^{\pm}(z_{1}),[e_{i}^{\pm}(z_{2}),e_{j}^{\pm}(w)]]=0\quad\textnormal{{ if }}a_{i,j}=-1,$
@@
This implies that relation (L3) holds on ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$. Note that in the presence of relations (L4) and (L5), (L6) is equivalent to the order-$2$ Serre relation. Then it follows from Remark 5.14 that ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$ is an $\widehat{\mathcal{L}}_{A}$-module of level $\ell$. It is clear that ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$ is generated by vector ${\bf 1}+\hbar{\mathcal{V}}_{A}(\ell)$. ∎

### v2

@@ 5.2. Universal vacuum $\widehat{\mathcal{DY}}(A)$-module ${\mathcal{V}}_{A}(\ell)$
(5.41) $\tilde{h}_{i}(z)\tilde{h}_{j}(w)-\tilde{h}_{j}(w)\tilde{h}_{i}(z)$
$\quad=[r_{i}a_{i,j}]_{q^{\partial_{w}}}[\ell]_{q^{\partial_{w}}}((z-w+\ell\hbar)^{-2}-(w-z+\ell\hbar)^{-2}),$
(5.42) $\tilde{h}_{i}(z)\tilde{e}_{j}(w)-\tilde{e}_{j}(w)\tilde{h}_{i}(z)$
$\quad=\tilde{e}_{j}(w)[r_{i}a_{i,j}]_{q^{\partial_{w}}}((z-w+\ell\hbar)^{-1}-(w-z+\ell\hbar)^{-1}),$
(5.43) $\tilde{h}_{i}(z)\tilde{f}_{j}(w)-\tilde{f}_{j}(w)\tilde{h}_{i}(z)$
$\quad=-\tilde{f}_{j}(w)[r_{i}a_{i,j}]_{q^{\partial_{w}}}((z-w+\ell\hbar)^{-1}-(w-z+\ell\hbar)^{-1}),$
(5.44) $(z-w-r_{i}a_{i,j}\hbar)\tilde{e}_{i}(z)\tilde{e}_{j}(w)=(z-w+r_{i}a_{i,j}\hbar)\tilde{e}_{j}(w)\tilde{e}_{i}(z),$
(5.45) $(z-w-r_{i}a_{i,j}\hbar)\tilde{f}_{i}(z)\tilde{f}_{j}(w)=(z-w+r_{i}a_{i,j}\hbar)\tilde{f}_{j}(w)\tilde{f}_{i}(z),$
(5.46) $(z-w+r_{i}a_{i,j}\hbar)\tilde{e}_{i}(z)\tilde{f}_{j}(w)=(z-w-r_{i}a_{i,j}\hbar)\tilde{f}_{j}(w)\tilde{e}_{i}(z)\quad\text{ if }i\neq j,$
(5.47) $(z-w)(z-w+2\ell\hbar)\!\big((z-w+2r_{i}\hbar)\tilde{e}_{i}(z)\tilde{f}_{i}(w)$
$\qquad-(z-w-2r_{i}\hbar)\tilde{f}_{i}(w)\tilde{e}_{i}(z)\big)\!=0,$
and the Serre-like relations for $(\tilde{e}_{i}(z),\tilde{e}_{j}(z),1-a_{i,j})$ and $(\tilde{f}_{i}(z),\tilde{f}_{j}(z),1-a_{i,j})$ for $i,j\in I$ with $a_{i,j}\leq 0$.
Note that for any $s,t\in{\mathbb{Z}}$, the coefficients of monomial $z^{s}w^{t}$ in all the relations above indeed lie in $\widetilde{\mathcal{U}}$$\hbar$]].
@@
(5.57) $(\hat{e}_{i})_{0}(\hat{f}_{j})-\frac{\delta_{i,j}}{2r_{i}\hbar}({\bf 1}-C_{i}),\ \ (\hat{e}_{i})_{n}(\hat{f}_{j})-\frac{\delta_{i,j}\ell}{r_{i}}(-2\ell\hbar)^{n-1}C_{i}\ \ \text{ for }i,j\in I,\ n\geq 1.$
@@
$Y(\check{e}_{i},z)Y(\check{f}_{j},w)\sim\left(\frac{w-z+r_{i}a_{i,j}\hbar}{w-z-r_{i}a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z).$
@@
(5.62) $Y(\check{e}_{i},z)Y(\check{f}_{j},w)-\!\left(\frac{w-z+r_{i}a_{i,j}\hbar}{w-z-r_{i}a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z)$
@@
(5.63) $(\check{e}_{i})_{0}(\check{f}_{j})=\frac{\delta_{i,j}}{2r_{i}\hbar}({\bf 1}-\mathfrak{c}_{i}),\ \ (\check{e}_{i})_{n}(\check{f}_{j})=\frac{\delta_{i,j}\ell}{r_{i}}(-2\ell\hbar)^{n-1}\mathfrak{c}_{i}$
@@
(5.64) $Y(\check{e}_{i},z)Y(\check{f}_{j},w)-\!\left(\frac{w-z+r_{i}a_{i,j}\hbar}{w-z-r_{i}a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z)$
@@
$[Y(\check{h}_{i},z),Y(\check{h}_{j},w)]=[r_{i}a_{i,j}]_{q^{\partial_{w}}}[\ell]_{q^{\partial_{w}}}\left((z-w+\ell\hbar)^{-2}-(w-z+\ell\hbar)^{-2}\right)\!,$
$[Y(\check{h}_{i},z),Y(\check{e}_{j}^{\pm},w)]=\pm Y(\check{e}_{j}^{\pm},w)[r_{i}a_{i,j}]_{q^{\partial_{w}}}\left((z-w+\ell\hbar)^{-1}+(w-z+\ell\hbar)^{-1}\right)\!,$
$(z-w-r_{i}a_{i,j}\hbar)Y(\check{e}_{i}^{\pm},z)Y(\check{e}_{j}^{\pm},w)=(z-w+r_{i}a_{i,j}\hbar)Y(\check{e}_{j}^{\pm},w)Y(\check{e}_{i}^{\pm},z),$
$Y(\check{e}_{i},z)Y(\check{f}_{j},w)-\left(\frac{w-z+r_{i}a_{i,j}\hbar}{w-z-r_{i}a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z)$
$\quad=\delta_{i,j}\frac{1}{2r_{i}\hbar}\!\left(\!z^{-1}\delta\!\left(\frac{w}{z}\right)-Y\!\left(q^{(1-\ell)\mathcal{D}}E^{-}(-\check{h}_{i},-2\hbar){\bf 1},w\right)z^{-1}\delta\!\left(\frac{w-2\ell\hbar}{z}\right)\!\right)\!,$
where $\mathcal{D}$ denotes the $D$-operator of ${\mathcal{V}}_{A}(\ell)$. In addition, for $i,j\in I$ with $a_{i,j}\leq 0$, the Serre-like relations for $(Y(\check{e}_{i}^{\pm},z),Y(\check{e}_{j}^{\pm},z),1-a_{i,j})$ hold. On the other hand, for any ${\mathcal{V}}_{A}(\ell)$-module $(W,Y_{W})$, all of these relations hold on $W$ by Propositions 3.26 and 3.22 and by Lemma 3.25. Furthermore, we have
@@
(L1) $[h_{i}(z),h_{j}(w)]=r_{i}a_{i,j}\frac{\partial}{\partial{w}}z^{-1}\delta\!\left(\frac{w}{z}\right){\bf k},$
(L2) $[h_{i}(z),e_{j}^{\pm}(w)]=\pm r_{i}a_{i,j}e_{j}^{\pm}(w)z^{-1}\delta\!\left(\frac{w}{z}\right)\!,$
(L3) $[e_{i}^{+}(z),e_{j}^{-}(w)]=\frac{\delta_{i,j}}{r_{i}}\left(\!h_{i}(w)z^{-1}\delta\!\left(\frac{w}{z}\right)+{\bf k}\frac{\partial}{\partial{w}}z^{-1}\delta\!\left(\frac{w}{z}\right)\right)\!,$
@@
(L5) $(z-w)[e_{i}^{\pm}(z),e_{j}^{\pm}(w)]=0\quad\textnormal{{if }}a_{i,j}<0,$
(L6) $[e_{i}^{\pm}(z_{1}),[e_{i}^{\pm}(z_{2}),\dots[e_{i}^{\pm}(z_{1-a_{i,j}}),e_{j}^{\pm}(w)]\cdots]]=0\quad\textnormal{{ if }}a_{i,j}<0,$
@@
This implies that relation (L3) holds on ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$. Note that in the presence of relations (L4) and (L5), (L6) is equivalent to the Serre-like relation. Then it follows from Remark 5.14 that ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$ is an $\widehat{\mathcal{L}}_{A}$-module of level $\ell$. It is clear that ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$ is generated by vector ${\bf 1}+\hbar{\mathcal{V}}_{A}(\ell)$. ∎